# Differential Contrastive Training for Gaze Estimation

### v2

@@ 4.4 Ablation Study
Table 4: Effectiveness of the Differential Contrastive Training on within-domain tasks.
@@
GazeTR+DCTrain 3.73 5.15 10.50
Table 5: Effectiveness of the Differential Contrastive Training on cross-domain tasks.
@@
4.4.2 Ablation Study of Proposed Modules
To investigate the effectiveness of each proposed component, we compare the performance of some degraded models on both within and cross-domain tasks. Based on the fully model, we remove one proposed component each time and get the results in Table  6 and in Table  7. The $1st.$ rows refer the baseline models. In within-domain evaluation, the baseline model consists of a CNN-Transformer-based Feature Extractor and an MLP-based regressor. In cross-domain evaluation, the Transformer part of the Feature Extractor is replaced by a 3-layer MLP. In within-domain tasks, no matter which component is invalidated, the performance decreases. This proves the importance of each of them. 1) By comparing the $2nd.$ rows with the $5th.$ rows (full model), it demonstrates that the Semantic Difference-aware branch can help the visual branch capture robust and pure gaze-related features via image-text alignment. 2) By comparing the $4th.$ rows with the $5th.$ rows, the DGR has been proven effective in reducing degrees of freedom and improving generalizability. 3) By comparing the $3rd.$ rows with the $5th.$ rows, AFU could bring improvements to within-domain tasks but lead a decline in performance for cross-domain ones. We analyze the possible reasons. CLIP is likely to encode scene contexts (e.g., image resolution, illumination, etc.) into its obtained appearance embeddings. For cross-domain tasks, the test samples own significantly different environments with training ones. Thus the AFU is more adept at capturing the scene information of the source domain, which may disturb the estimation of disjoint target samples.
Table 6: Ablation study results of proposed components on within-domain tasks.
@@ 4.4.2 Ablation Study of Proposed Modules
Table 7: Ablation study results of proposed components on cross-domain tasks.
@@
Table 8: Analysis of differential gaze prompts. The fixed template is ‘The directions of gaze in the two photos are {Grade Name}.’
@@
4.4.3 Analysis of Differential Gaze Prompts
In this section, we evaluate our DCGaze under different prompt designs (Table  8). Firstly, we vary the number of grade prompts $K$ from 2 to 5. The results show that increasing the number of grades can better help the model to identify the subtle gaze differences between facial images, leading to better feature representations. However, the more levels there are, the more difficult manual designs of textual prompts become. Thus, we select 5 grades in our experiments. Secondly, we explore different strategies to describe these prompts. To be specific, we use different words for each differential grade name or propose a learnable prompt template following the CoOp method [54]. Two main observations could be concluded from the comparisons. 1) The manual prompts combining with degree adverbs can convey more precise semantic information, which lead to better results. 2) The learnable prompt template performs worse than the fixed one, whose possible reason is that the fixed prompt template could clearly describe the gaze difference between images while the learnable one may introduce some noise during the learning process.
4.4.4 Analysis of Feature Refinement Strategy
@@
$\centering\begin{split}Q_{pry}=f^{pry}W_{Q},K_{clip}=&f^{clip}W_{K},{V}_{clip}=f^{clip}W_{V},\\ f^{img}=\text{Softmax}&(Q_{pry}K_{clip}^{T}/\beta){V}_{clip}.\end{split}\@add@centering$ (11)
$\bullet$ ‘Gated Information Fusion’: Following [1], the prior appearance feature $f^{clip}$ is processed by a sigmoid operation to generate a mask, which is used to activate the primary appearance feature $f^{img}$ via Hadamard product. The detailed process refers to Eq. (12):
$f^{img}=\ f^{pry}\circ\text{Sigmoid}(f^{clip}),$ (12)
@@
Table 9: Analysis of feature fusion strategy.
@@
4.5 More Discussion  [^1] In supplementary materials, we conduct more experiments to evaluate the properties of our proposed modules.
4.5.1 Visualization of Obtained Gaze Features
To quantitatively demonstrate the advantages of our enhanced gaze features, we visualize the distribution of the training samples of Gaze360 and ETH-XGaze datasets by t-SNE [35], following [41]. We select all the training samples of Gaze360 dataset and randomly select 10000 training samples from ETH-XGaze dataset. In the scatter plot, the samples are clustered by KMeans [21] according to their gaze labels, so that the samples with similar gaze directions share similar colors. The feature distributions of the Baseline and our DCGaze are shown in Figure 3. As shown in the left figure, the sample points of baseline are scattered in a chaotic manner. By contrast, feature points of our enhanced features are distributed in an organized way, in which the features with similar gaze directions are clustered and can be regressed to similar gazes. It demonstrates our enhanced features being purified gaze-related ones, which effectively improve the discriminability and generalization.
Figure 3: Visualization of obtained gaze features of Gaze360 (upper) and ETH-XGaze (bottom) datasets.
4.5.2 Visualization of Estimated Gazes
@@
Figure 4: Visualization of estimated gazes.
5 Conclusion
@@
In this paper, we have proposed a novel Differential Contrastive Gaze Estimation network (DCGaze) for gaze estimation, which leverages the powerful capabilities of the pre-trained CLIP to improve the representation capacity of a primary network by a novel Differential Contrastive Training strategy. Firstly, in the Visual Appearance-aware branch, the AFU and DGR have been proposed to help the primary gaze estimation model to capture informative and gaze-related features. Besides, to realize the Differential Contrastive Training, the Semantic Difference-aware branch has constructed image-pairs and designed differential gaze prompt for each of them. Through visual-semantic alignment, the primary gaze estimation network has been endowed with the ability of extracting robust and pure gaze-related features. In extensive experiments, DCGaze has achieved remarkable performance on both within-domain tasks and cross-domain tasks. Limitation: DCGaze is not a lightweight model, which would limit its deployment on edge devices in real application scenarios. Therefore, in the future, we would keep improving DCGaze in terms of reducing its scalability while maintaining the precision of prediction.  [^3] In supplementary materials, we discuss the scalability of DCGaze compared to other methods.

### v3

@@ 4.4. Ablation Study
Table 4. Effectiveness of the DCTrain on within-domain tasks.
@@
GazeTR+DCTrain 3.73 5.15 10.56
Table 5. Effectiveness of the Differential Contrastive Training on cross-domain tasks (SDA).
@@
4.4.2. Ablation Study of Proposed Modules
[…]
Table 6. Ablation study results of proposed components on within-domain tasks.
@@ 4.4.2. Ablation Study of Proposed Modules
Table 7. Ablation study results of proposed components on cross-domain tasks (SDA).
@@
Table 8. Analysis of differential gaze prompts. The fixed template is ‘The directions of gaze in the two photos are {Grade Name}.’
@@
4.4.3. Analysis of Differential Gaze Prompts
In this section, we evaluate our DCGaze under different prompt designs (Table  8). Firstly, we vary the number of grade prompts $K$ from 2 to 5. The results show that increasing the number of grades can better help the model to identify the subtle gaze differences between facial images, leading to better feature representations. However, the more levels there are, the more difficult manual designs of textual prompts become. Thus, we select 5 grades in our experiments. Secondly, we explore different strategies to describe these prompts. To be specific, we use different words for each differential grade name or propose a learnable prompt template following the CoOp method (Zhou et al., 2022). Two main observations could be concluded from the comparisons. 1) The manual prompts combining with degree adverbs can convey more precise semantic information, which lead to better results. 2) The learnable prompt template performs worse than the fixed one, whose possible reason is that the fixed prompt template could clearly describe the gaze difference between images while the learnable one may introduce some noise during the learning process.
4.4.4. Analysis of Feature Refinement Strategy
@@
(11) $\centering\begin{split}Q_{pry}=f^{pry}W_{Q},K_{clip}=&f^{clip}W_{K},{V}_{clip}=f^{clip}W_{V},\\ f^{img}=\text{Softmax}&(Q_{pry}K_{clip}^{T}/\beta){V}_{clip}.\end{split}\@add@centering$
$\bullet$ ‘Gated Information Fusion’: Following (Arevalo et al., 2017), the prior appearance feature $f^{clip}$ is processed by a sigmoid operation to generate a mask, which is used to activate the primary appearance feature $f^{img}$ via Hadamard product. The detailed process refers to Eq. (12):
(12) $f^{img}=\ f^{pry}\circ\text{Sigmoid}(f^{clip}),$
@@
Table 9. Analysis of feature fusion strategy.
@@
4.5. More Discussion  [^1] In supplementary materials, we conduct more experiments to evaluate the properties of our proposed modules.
4.5.1. Visualization of Obtained Gaze Features
To quantitatively demonstrate the advantages of our enhanced gaze features, we visualize the distribution of the training samples of Gaze360 and ETH-XGaze datasets by t-SNE (van der Maaten and Hinton, 2008), following (Wang et al., 2022a). We select all the training samples of Gaze360 dataset and randomly select 10000 training samples from ETH-XGaze dataset. In the scatter plot, the samples are clustered by KMeans (Ikotun et al., 2023) according to their gaze labels, so that the samples with similar gaze directions share similar colors.
Figure 3. Visualization of obtained gaze features of Gaze360 (upper) and ETH-XGaze (bottom) datasets.
The feature distributions of the Baseline and our DCGaze are shown in Figure 3. As shown in the left figure, the sample points of baseline are scattered in a chaotic manner. By contrast, feature points of our enhanced features are distributed in an organized way, in which the features with similar gaze directions are clustered and can be regressed to similar gazes. It demonstrates our enhanced features being purified gaze-related ones, which effectively improve the discriminability and generalization.
4.5.2. Visualization of Estimated Gazes
@@
Figure 4. Visualization of estimated gazes.
5. Conclusion
@@
In this paper, we have proposed a novel Differential Contrastive Gaze Estimation network (DCGaze) for gaze estimation, which leverages the powerful capabilities of the pre-trained CLIP to improve the representation capacity of a primary network by a novel Differential Contrastive Training strategy. Firstly, in the Visual Appearance-aware branch, the AFU and DGR have been proposed to help the primary gaze estimation model to capture informative and gaze-related features. Besides, to realize the Differential Contrastive Training, the Semantic Difference-aware branch has constructed image pairs and designed differential gaze prompt for each of them. Through visual-semantic alignment, the primary gaze estimation network has been endowed with the ability of extracting robust and pure gaze-related features. In extensive experiments, DCGaze has achieved remarkable performance on both within-domain tasks and cross-domain tasks. Limitation: DCGaze is not a lightweight model, which would limit its deployment on edge devices in real application scenarios. Therefore, in the future, we would keep improving DCGaze in terms of reducing its scalability while maintaining the precision of prediction.  [^3] In supplementary materials, we discuss the scalability of DCGaze compared to other methods.
Acknowledgements.
This work is supported by Beijing Natural Science Foundation L251032, 4242028, the National Natural Science Foundation of China 62376021, Natural Science Foundation of Hebei Province F2025105036.
Appendix A Details of the early attempt in Introduction
As discussed in Introduction, to explore this concern ‘can CLIP understand human gaze?’, we make a simple early attempt. The pipeline of our conducted experiment is shown in Figure S-5.
[…]
(S-1) $f_{i}^{clip}=ImageEncoder\left(x_{i}\right),$
(S-2) $f_{j}^{text}=TextEncoder\left(t_{j}\right),j=1,2,3,4,$
(S-3) $S_{ij}=\frac{f_{i}^{clip}\cdot{{(f}_{j}^{text})}^{T}}{||f_{i}^{clip}||\ {||f}_{j}^{text}||},$
(S-4) $g_{i}=\sum_{j=1}^{4}{S_{ij}\times}g_{j}^{bin}.$
Obviously, the above simple CLIP-based gaze estimation method does not require any training process, which only relies on the pre-trained CLIP model. We evaluate the performance of this simple attempt on MPIIFaceGaze dataset, whose average angular error is reported in Table S-2.
For comparison purposes, we train a traditional ResNet-based (including ResNet18 and ResNet50) gaze estimation network on ETH-XGaze dataset and test it on MPIIFaceGaze dataset. To be specific, the traditional network takes a facial image as input, extracts a gaze feature of it via several convolutional layers and gets a final gaze via a fully connected layer. From the results we find that even without training on any exclusive gaze estimation datasets, the CLIP achieves performance comparable to ResNet18 and ResNet50 that trained with cross-domain samples. Therefore, we draw a conclusion that the CLIP could understand human gazes subtly, but its potentials have not been fully exploited.
Table S-1. The correspondences among prototypes, gaze bins and prompts.
Prototype
Gaze Bin
[…]
Prompt
up [0, $\frac{\pi}{2}$] A photo of a face gazing up.
down [0, -$\frac{\pi}{2}$] A photo of a face gazing down.
left [$\frac{\pi}{2}$, 0] A photo of a face gazing left.
right [-$\frac{\pi}{2}$, 0] A photo of a face gazing right.
In this paper, we develop the potentials of the pre-trained CLIP in boosting gaze estimation performance from a real new perspective, which puts CLIP into a supporting role to facilitate the extraction of gaze-related features of a primary gaze estimation network.
Figure S-5. The pipeline of our early attempt.
Table S-2. The results of ResNet and CLIP on MPIIFaceGaze dataset. We evaulate ResNet18 and ResNet50 on the cross-domain task $D_{E}\rightarrow D_{M}$ which indicates training on ETH-XGaze and testing on MPIIFaceGaze dataset.
Method Angular error on MPIIFaceGaze
ResNet18 ($\mathcal{D}_{E}\rightarrow\mathcal{D}_{M}$) $9.40^{\circ}$
ResNet50 ($\mathcal{D}_{E}\rightarrow\mathcal{D}_{M}$) $12.71^{\circ}$
CLIP $11.10^{\circ}$
Figure S-6. The results of different groups of $\alpha$ and $\beta$ on MPIIFaceGaze, EyeDiap and Gaze360 datasets.
Appendix B Introduction of the datasets used in our experiments
Our method is evaluated on four popular gaze estimation datasets, which are MPIIFaceGaze, EyeDiap, Gaze360 and ETH-XGaze.
MPIIFaceGaze (Zhang et al., 2019) dataset is captured by web cameras in daily life. It contains 45k images from 15 subjects.
EyeDiap (Funes Mora et al., 2014) dataset is collected in laboratory environment with screen targets or floating targets. It contains 16k images of the video clips from 14 subjects.
Gaze360 (Kellnhofer et al., 2019) dataset is captured in both indoor and outdoor environments via a 360° camera. It contains 101k images from 238 subjects with a wide-range head pose and gaze directions.
ETH-XGaze (Zhang et al., 2020) dataset is collected by high resolution cameras in laboratory environment. It contains 1.1M images from 110 subjects. For our experiments, we select 756,540 images from 80 subjects among them as the training set.
Appendix C More Discussion
C.1. Analyses of hyperparameter
As discussed in Optimization and Inference, there are three constraints of our DCGaze, which are $L_{gaze}$, $L_{mask}$ and $L_{align}$. Their weights are 1, $\alpha$ and $\beta$, respectively. In this section, we discuss the influence of $\alpha$ and $\beta$ in detail. We fix $\alpha$ ($\alpha=0.1$) and vary $\beta$ in the range of [0, 0.01, 0.1, 1]. Inversely, we fix $\beta$ ($\beta=0.1$) and vary $\alpha$ in the range [0, 0.01, 0.1, 1]. The experiments are conducted on MPIIFaceGaze, EyeDiap and Gaze360 datasets over the within-domain tasks. As the results shown in Figure S-6, there are little variation of angle errors when the hyperparameters $\alpha$, $\beta$ are set to different selected values, which means the DCGaze is insensitive to those hyperparameters.
C.2. More visualization results
(a) Visualization of features of 5 subjects from MPIIFaceGaze dataset.
(b) Visualization of features of 5 subjects from EyeDiap dataset.
Figure S-7. Visualization of obtained gaze features on MPIIFaceGaze and EyeDiap datasets, where the feature points belonging to the same person share the same color.
As illustrated in More Discussion, we visualize the distribution of the enhanced features of Gaze360 and ETH-XGaze datasets. In this section, we add more visualization results to demonstrate the advantages of our enhanced gaze features.
We visualize the distributions of gaze features from a subject-specific perspective, in which the feature points belonging to the same person share the same color. We randomly select 5 test subjects from the $D_{E}\rightarrow D_{M}$ and $D_{E}\rightarrow D_{D}$ tasks and visualize all the gaze features of them obtained via the Baseline and the DCGaze, respectively. Those features are also visualized in 2D using t-SNE (van der Maaten and Hinton, 2008). The visualization results are shown in Figure S-7. It is evident that the features captured by Baseline of each individual are gathered together because they still contain some identity information which is unrelated to gaze. Compared to baseline method, our enhanced features of the same subject are more dispersed, which proves that our DCGaze effectively reduces the variations among subjects and learns purified gaze-related representations.
C.3. Discussion of scalability of DCGaze
The number of trainable parameters of our DCGaze and other existing gaze estimation methods are shown in Table S-3 and Table S-4 . The scalability of comparison methods is reported in their original papers or counted via reproducing their released codes. In terms of the number of parameters, our model is comparable to the several lightweight models. Therefore, in the future, we would keep improving DCGaze via reducing its scalability while maintaining the precision of prediction.
Table S-3. Comparisons of scalability with existing methods on within-domain tasks.
Model Gaze error Number of
MPII EyeDiap Gaze360 parameters
RT-GENE (Fischer et al., 2018) 4.30 5.90 - 82.0M
CA-Net (Cheng et al., 2020a) 4.37 5.27 11.20 34.1M
PureGaze (Cheng et al., 2022) 4.13 5.82 - 31.55M
GazeTR-Pure (Cheng and Lu, 2022) 4.74 5.72 - 227.3M
DCGaze 3.71 4.97 10.54 34.99M
Table S-4. Comparisons of scalability with existing methods on cross domain tasks.
Model Gaze error
Number of
Parameters
$\mathcal{D}_{E}$
$\rightarrow\mathcal{D}_{M}$
$\mathcal{D}_{E}$
$\rightarrow\mathcal{D}_{D}$
$\mathcal{D}_{G}$
$\rightarrow\mathcal{D}_{M}$
$\mathcal{D}_{G}$
$\rightarrow\mathcal{D}_{D}$
Gaze360 (Kellnhofer et al., 2019) 6.24 7.47 7.17 7.66 14.6M
PnP-GA (Liu et al., 2021) 5.53 5.87 6.18 7.92 116.9M
DCGaze 6.47 6.70 6.45 7.74 34.99M